# Primal Methods for Variational Inequality Problems with Functional Constraints

### v1

@@ 4.3 The Special Case with Convex Minimization
This is equivalent to the variational inequality problem associated with the operator $F(x):=\nabla f(x)$ and the feasible set $\mathcal{C}=\{x\in\mathbb{R}^{d}|g_{i}(x)\leq 0,\forall i\in[m]\}$. The guarantees of CGM in this setting are as follows.
@@
Suppose $f(x)$ is $\mu$-strongly-convex, $\ell_{f}$-smooth, and $L_{f}$-Lipschitz, and $g_{i}(x)$ is convex, $\ell_{g}$-smooth, and $L_{g}$-Lipschitz for each $i\in[m]$ such that the feasible set $\mathcal{C}$ is contained in an Euclidean ball of radius $D$. Let $T$ be large enough such that $T\geq\max\{3,\kappa_{f}\}(\log T)$, where $\kappa_{f}:=\ell_{f}/\mu$ denotes the condition number. Then CGM with stepsize $\eta=(\log T)/(\mu T)$ and $\alpha=\mu$ satisfies
@@
since each $g_{i}(x)$ is convex and $\alpha>0$. This suggests that $\alpha(x-x_{t})\in V_{\alpha}(x_{t})$ for any $x\in\mathcal{C}$. By the computation of $v_{t}$, we then have that $\lVert v_{t}+\nabla f(x_{t})\rVert^{2}\leq\lVert\alpha(x^{*}-x_{t})+\nabla f(x_{t})\rVert^{2}$ for $x^{*}=\arg\min_{x\in\mathcal{C}}f(x)$. Using smoothness of $f(x)$, we conclude
@@
where we use the assumption that $f(x)$ is $\mu$-strongly-convex. When setting $\alpha\leq\mu$, we obtain
@@ 6.2 Matrix Game with Quadratic Constraints
where $z=(x,y)\in\mathbb{R}^{2d}$ is the concatenation of $x$ and $y$. The problem is equivalent to the constrained variational inequality problem associated with $F(z)=(Ay,-A^{\top}(x-a))$ and $\mathcal{C}=\{z\in\mathbb{R}^{2d}\,|\,(1/2)z^{\top}Bz\leq c\}$. Here, $A\in\mathbb{R}^{d\times d}$ is a random matrix with each entry sampled from $\mathcal{N}(0,1)$, $a\in\mathbb{R}^{d}$ is a vector with each entry sampled from $\mathcal{N}(0,0.1)$, $B\in\mathbb{R}^{2d\times 2d}$ is a random positive definite matrix with each eigenvalue uniformly sampled from $[0.1,10]$, and $c>0$ is sampled uniformly from $[0.1,10]$. For any point $\hat{z}=(\hat{x},\hat{y})$, its optimality is measured by the (strong) gap $\max_{z\in\mathcal{C}}F(\hat{z})^{\top}(\hat{z}-z)=-\min_{z\in\mathcal{C}}z^{\top}F(\hat{z})=(2c\,F(\hat{z})^{\top}B^{-1}F(\hat{z}))^{1/2}$ through solving the corresponding KKT system, and its feasibility is measured by the constraint violation $\max\{0,(1/2)\hat{z}^{\top}B\hat{z}-c\}$.
@@ 7 Conclusion
We present a primal method for solving monotone variational inequality problems with general functional constraints. Our algorithm achieves the same complexity on querying $F$ as projection-based methods and does not rely on the optimal Lagrange multipliers unlike existing primal-dual methods. Several interesting questions have yet to be explored. The current analysis of CGM (Algorithm 1) for the strongly-monotone setting achieves a rate slightly worse than $\mathcal{O}(1/T)$ on the constraint violation, where $T$ is the number of iterations. It remains unknown whether such guarantees can be improved. An exploration into the monotone and Lipschitz case is also valuable, particularly to determine if the improved complexity $\mathcal{O}(1/\epsilon)$ can be achieved, as observed in the context of projection-based methods, such as extragradient and optimistic gradient [41, 40]. Furthermore, extensions to stochastic settings [28, 1] and non-monotone settings [65, 12] are left for future investigations.
@@ Appendix A Proof of Lemma 1
We recall that $V_{\alpha}(x_{t})=\{v\in\mathbb{R}^{d}\,|\,\alpha g_{i}(x_{t})+\nabla g_{i}(x_{t})^{\top}v\leq 0,\forall i\in I_{x_{t}}\}$. The equation above suggests that $\alpha(x-x_{t})\in V_{\alpha}(x_{t})$ for any $x\in\mathcal{C}$. By the guarantee that $(v_{t}+F(x_{t}))^{\top}(v_{t}-v)\leq\epsilon/2$, $\forall v\in V_{\alpha}(x_{t})$, we know that for any $v\in V_{\alpha}(x_{t})$,
@@
Since $\gamma>1$, by basic arithmetic calculations, this suggests that
@@
$=\Big{(}1-\alpha\eta_{k}+(1-\gamma^{-1})\alpha\eta_{k}\Big{)}(\lVert x_{k}\rVert^{2}-D^{2})$
$\qquad+(1-\gamma^{-1})\alpha\eta_{k}D^{2}+(\gamma-1)\alpha\eta_{k}\left(D+\frac{4L_{F}}{\alpha}\right)^{2}$
@@ Appendix B CGM with Simplex Constraints
This suggests that the solution of $\lambda$ is
@@
$=\frac{d-n}{d-n+1}(q_{1}+\lambda)$
@@
This suggests that $q_{i}+\lambda\leq 0,\forall i\in N$ and thus $q=0$.
@@
$q_{j}+\frac{1}{d-n+j}\left(1-s_{\bar{N}}-\sum_{i=1}^{j}q_{i}\right)$ $=\frac{1}{d-n+j}\left(1-s_{\bar{N}}-\sum_{i=1}^{\rho}q_{i}+(d-m+j)q_{j}-\sum_{i=\rho+1}^{j}q_{i}\right)$
@@
This suggests that $\rho=\max J$ and leads to the method described in Algorithm 3. Since the algorithm only requires to sort $m$ coordinates instead of $d$, it is provably more efficient compared to the projection oracle. The difference is particularly pronounced if the quadratic program is sparse.

### v2

@@ 4.3 The Special Case with Convex Minimization
This is equivalent to the variational inequality problem associated with the operator $F(x):=\nabla f(x)$ and the feasible set $\mathcal{C}=\{x\in\mathbb{R}^{d}|g_{i}(x)\leq 0,\forall i\in[m]\}$. The guarantees of CGM in this setting are as follows. Note that the guarantees in this case apply to the last iterate $x_{T}$.
@@
Suppose $f(x)$ is $\mu$–strongly-convex, $\ell_{f}$-smooth, and $L_{f}$-Lipschitz, and $g_{i}(x)$ is convex, $\ell_{g}$-smooth, and $L_{g}$-Lipschitz for each $i\in[m]$ such that the feasible set $\mathcal{C}$ is contained in an Euclidean ball of radius $D$. Let $T$ be large enough such that $T\geq\max\{3,\kappa_{f}\}(\log T)$, where $\kappa_{f}:=\ell_{f}/\mu$ denotes the condition number. Then CGM with $\epsilon=0$ for simplicity, stepsize $\eta=(\log T)/(\mu T)$, and $\alpha=\mu$ satisfies that
@@
since each $g_{i}(x)$ is convex and $\alpha>0$. This implies that $\alpha(x-x_{t})\in V_{\alpha}(x_{t})$ for any $x\in\mathcal{C}$. By the computation of $v_{t}$, we then have that $\lVert v_{t}+\nabla f(x_{t})\rVert^{2}\leq\lVert\alpha(x^{*}-x_{t})+\nabla f(x_{t})\rVert^{2}$ for $x^{*}=\arg\min_{x\in\mathcal{C}}f(x)$. Using smoothness of $f(x)$, we conclude
@@
where we use the assumption that $f(x)$ is $\mu$–strongly-convex. When setting $\alpha\leq\mu$, we obtain
@@ 6.2 Matrix Game with Quadratic Constraints
where $z=(x,y)\in\mathbb{R}^{2d}$ is the concatenation of $x$ and $y$. The problem is equivalent to the constrained variational inequality problem associated with $F(z)=(Ay,-A^{\top}(x-a))$ and $\mathcal{C}=\{z\in\mathbb{R}^{2d}\,|\,(1/2)z^{\top}Bz\leq c\}$. Here, $A\in\mathbb{R}^{d\times d}$ is a random matrix with each entry sampled from $\mathcal{N}(0,1)$, $a\in\mathbb{R}^{d}$ is a vector with each entry sampled from $\mathcal{N}(0,0.1)$, $B\in\mathbb{R}^{2d\times 2d}$ is a random positive definite matrix with each eigenvalue uniformly sampled from $[0.1,10]$, and $c>0$ is sampled uniformly from $[0.1,10]$. For any point $\hat{z}=(\hat{x},\hat{y})$, its optimality is measured by the (strong) gap $\max_{z\in\mathcal{C}}F(\hat{z})^{\top}(\hat{z}-z)$, effectively the same as $-\min_{z\in\mathcal{C}}z^{\top}F(\hat{z})=(2c\,F(\hat{z})^{\top}B^{-1}F(\hat{z}))^{1/2}$ through solving the corresponding KKT system, and its feasibility is measured by the constraint violation $\max\{0,(1/2)\hat{z}^{\top}B\hat{z}-c\}$.
@@ 7 Conclusion
We present a primal method for solving monotone variational inequality problems with general functional constraints. Our algorithm achieves the same complexity on querying $F$ as projection-based methods and does not rely on any information about the optimal Lagrange multipliers unlike existing primal-dual methods. Several interesting questions have yet to be explored. The current analysis of CGM (Algorithm 1) for the strongly-monotone setting achieves a rate slightly worse than $\mathcal{O}(1/T)$ on the constraint violation, where $T$ is the number of iterations. It remains unknown whether such guarantees can be improved. Recently, Zamani and Glineur [95] presented a refined analysis of the projected subgradient method, demonstrating last iterate convergence for nonsmooth convex optimization. Extending our analysis in this direction could be interesting. An exploration into the monotone and Lipschitz case is also valuable, particularly to determine if the improved complexity $\mathcal{O}(1/\epsilon)$ can be achieved, as observed in the context of projection-based methods, such as extragradient and optimistic gradient [60, 59]. Furthermore, extensions to stochastic settings [41, 1, 49] and non-monotone settings [91, 20] are left for future investigations.
@@ Appendix A Proof of Lemma 1
We recall that $V_{\alpha}(x_{t})=\{v\in\mathbb{R}^{d}\,|\,\alpha g_{i}(x_{t})+\nabla g_{i}(x_{t})^{\top}v\leq 0,\forall i\in I_{x_{t}}\}$. The equation above implies that $\alpha(x-x_{t})\in V_{\alpha}(x_{t})$ for any $x\in\mathcal{C}$. By the guarantee that $(v_{t}+F(x_{t}))^{\top}(v_{t}-v)\leq\epsilon/2$, $\forall v\in V_{\alpha}(x_{t})$, we know that for any $v\in V_{\alpha}(x_{t})$,
@@
Since $\gamma>1$, by basic arithmetic calculations, this implies that
@@
$=\Big{(}1-\alpha\eta_{k}+(1-\gamma^{-1})\alpha\eta_{k}\Big{)}(\lVert x_{k}\rVert^{2}-D^{2})+(1-\gamma^{-1})\alpha\eta_{k}D^{2}+(\gamma-1)\alpha\eta_{k}\left(D+\frac{4L_{F}}{\alpha}\right)^{2}$
@@ Appendix B CGM with Simplex Constraints
This implies that the solution of $\lambda$ is
@@
$=\frac{d-n}{d-n+1}(q_{1}+\lambda)+\frac{1}{d-n+1}\sum_{i=1}^{\rho}(q_{i}+\lambda)$
@@
This implies that $q_{i}+\lambda\leq 0,\forall i\in N$ and $\rho=0$.
@@
$q_{j}+\frac{1}{d-n+j}\left(1-s_{\bar{N}}-\sum_{i=1}^{j}q_{i}\right)$ $=\frac{1}{d-n+j}\left(1-s_{\bar{N}}-\sum_{i=1}^{\rho}q_{i}+(d-n+j)q_{j}-\sum_{i=\rho+1}^{j}q_{i}\right)$
@@
This implies that $\rho=\max J$ and leads to the method described in Algorithm 3. Since the algorithm only requires to sort $n$ coordinates instead of $d$, it is provably more efficient compared to the projection oracle. The difference is particularly pronounced if the quadratic program is sparse.